There are five cars on display as prizes, and their five ignition keys are in box, You get to pick one key out of the box and try it in the ignition of one car. If it fits, you win the car. What are your chances of winning a car?

Suppose if you are in the audience, find the probability that the contestant will win the car you have selected.
Answer #1

1.

Sample space (total number of combinations between keys and cars) = 5 keys x 5 cars = 25 possible situations

Favorable cases = 5 (one of the 5 cars can be won)

Probability = Favorable cases / Sample space = 5/25 = 0.2

2. If a car is already chosen by me, only 1 of the 5 keys to choose from will fit the selected car. Essentially there is only 1 car and 5 keys, one of which is the correct key. So probability is 1/5 or 0.2